# Effective Hamiltonian for Twisted TMDs which Exhibit Pressure Dependent Topological Phase Transitions

### v1

@@ Abstract
Motivated by recent studies on topologically nontrivial moiré bands in twisted bilayer transition metal dichalcogenides (TMDs), we study twisted MoTe2 bilayer systems subject to pressure, which is applied perpendicular to the material surface. We start our investigation by first considering untwisted bilayer systems with an arbitrary relative shift between layers; a symmetry analysis for this case permits us to obtain a simple, effective, low-energy Hamiltonian. Ab initio density functional theory (DFT) was then employed to obtain relaxed geometric structures for pressures within the range of 0.0 - 3.5 GPa and corresponding band structures. The DFT data was then fitted to the low-energy Hamiltonian to obtain a pressure-dependent Hamiltonian. We then could model a twisted system by treating the twist as a position-dependent shift between layers. In summary, this approach allowed us to obtain the explicit analytical expressions for a Hamiltonian that describes a twisted MoTe2 bilayer under pressure. Our Hamiltonian then permitted us to study the impact of pressure on the band topology of the twisted system. As a result, we identified many pressure-induced topological phase transitions as indicated by changes in valley Chern numbers. Moreover, we found that pressure could be employed to flatten bands in some of the cases we considered.
@@ I Introduction
Exotic phenomena have also been studied in twisted TMDs (tTMDs). Experimental works, for instance, report the observation of Mott insulating state [19, 20] and quantum anomalous Hall effect [21]. On the theoretical side, twisted MoTe2 homobilayers have been predicted to exhibit nontrivial topological moiré bands [22], which indicates the existence of Chern insulators in this material. This prediction was followed by experimental results of several groups [23, 24, 25, 26, 27, 28], which confirmed signatures of the theoretical predictions in their data. One might also add another TMD layer to obtain so-called trilayer TMDs. Trilayer TMDs leave additional room for modification and have been shown to give rise to a plethora of Chern transitions[29]. Twisted TMDs also act as an exciting playground for simulating idealized models such as Hubbard [30, 31] and generalized Kane-Mele models [22, 32]. In some recent work [33], authors managed to construct an interacting model Hamiltonian. In the same work Ref. [33], they considered different twist angles and managed to predict transitions between various phases, including quantum anomalous Hall insulator (QAHI), quantum spin Hall insulator (QSHI), and various antiferromagnetic phases.
@@
This additional knob could be chosen in various ways. For instance, it could be circularly polarized light, which has been studied for twisted bilayer graphene [34, 35, 36, 37, 38] and TMDs [39]. Another additional knob that can be introduced to the moiré system is pressure. There are tremendous efforts underway, both experimentally [40, 41, 42, 43] and theoretically [44, 45, 46, 47], to study the effect of pressure on graphene-based moiré systems to employ it to tune its physical properties. As for TMDs, Ref. [48] studied the effect of pressure on the stability of the fractional Chern insulator phase in twisted WSe2. Other research groups also try to introduce electric fields [49] and strains [50, 51, 52, 53] to tune further the properties of graphene-based moiré materials.
@@
In this paper, we want to help simplify the study of tTMDs subjected to pressure. We present an explicit pressure-dependent low-energy Hamiltonian for a twisted MoTe2 homobilayer in the non-interacting limit to achieve this goal. The model was derived from a general symmetry-based guess [54]. First principle density functional theory (DFT) [55, 56] was employed to obtain band structures under pressure. The energy around a small region of $\mathbf{K}$ valley, obtained from the DFT computation at each pressure, was then fitted to the low-energy Hamiltonian from our symmetry analysis to obtain explicit expression of pressure-dependent coupling parameters. We also limit the range of pressures in which our model is valid. Further investigation then lead us to the prediction of pressure-induced topological phase transitions as indicated by jumps in valley Chern numbers that correspond to specific energy bands.
@@
The remainder of the paper is structured as follows. In Section II, we present the low-energy Hamiltonian of the twisted TMD and explain the symmetry-based methods that were employed to obtain it. We demonstrate the steps to obtain the coupling parameters by fitting them to DFT data and explain the methods to compute valley Chern numbers and energy bands. Then, in Section III, we show how we use DFT computations to study structures under pressure and obtain their energy bands. Subsequently, we use this data to find the expression of the pressure-dependent coupling parameters of the Hamiltonian. We further investigate the physics of the pressurized twisted TMD by computing valley Chern number of particular energy bands within the range of 0 - 3.5 GPa of pressure for twist angle $\theta$ of $1.4^{\circ}$ and $2.3^{\circ}$ and observe topological transitions that occur in our pressure range. Lastly, we conclude by summarizing our main findings in Section IV.
@@ II Material without pressure
This study focuses on MoTe2 homobilayer systems -including twisted bilayers. Fig. 1 illustrates such systems for the untwisted case - specifically, the MoTe2 bilayer in the AA and AB/BA stacking configurations. Before we study the twisted bilayer system, we first discuss the untwisted bilayer case to gain some physical understanding.
@@
To understand the geometry of layer shifts, we consider a so-called AA stacked MoTe2 bilayer system first (Fig. 1(a) –here, atomic positions of atoms in the $x-y$ plane of one layer are exactly above thos in the second layer). We may introduce a displacement vector
@@
which allows the top layer to be shifted with respect to the bottom layer. Choosing $n=\pm 1$ corresponds to the so-called AB/BA stacking (Fig. 1(b)) while $n=0$ corresponds to the AA stacking we used as starting point.
@@
We allowed the system to fully relax at a pressure of 0.0 GPa to obtain accurate lattice constants. With this approach, we computed the following lattice constants $a_{0}$ (see Eq. (1) and Fig. 1). For the unshifted case ($\mathbf{d}_{0}$) we found $a_{0}=3.482$ while for the shifted case($\mathbf{d}_{\pm 1}$) $a_{0}=3.487$. These numbers are very close to the lattice constants obtained for monolayer MoTe2 using DFT calculation with the LDA [58]. They differ only slightly due to coupling between the layers. Additionally, the vertical distance between the Mo atoms of each layer ($D_{z}$) was found to be 7.79Å for the unshifted structure and 6.91Å for the shifted structures, which also agrees with Ref. [22].
@@
It is clear from Fig. 2 that the two topmost energy bands near the so-called $\mathbf{K}$-point behave approximately quadratically. It is worth noting that, like other TMD materials, MoTe2 features a strong spin-orbit coupling that causes the splitting of the top valence bands. Particularly, they correspond to spin-up ($\uparrow$) states for the $K$ valley (spin-down ($\downarrow$) states for $K^{\prime}$ valley, which is due to the time-reversal symmetry) [59, 22]. We are now interested in finding an effective Hamiltonian that describes these two bands near $K$ for arbitrary shifts between layers. In the cases we consider, we find that one can write a two-band model Hamiltonian as
@@
which acts on the wave function of the top and bottom layers $(\phi_{t},\phi_{b})$. (Note: One may confirm that the upper band near $K$ has contributions mainly from the top layer, and the lower band has contributions mainly from the bottom layer, which suggests these identifications.)
@@ II.1.1 Intra-layer energies and their shift dependence
Figure 3: The reciprocal lattice vectors $\mathbf{G}_{j}$ connect point $K$ to the nearest equivalent $K$ points.
Here, $\mathbf{G}_{j}$ are the reciprocal lattice vectors that connect the origin, which is the $\mathbf{K}$ valley, to the other equivalent $\mathbf{K}$ valleys. We only choose the lowest-harmonics $\mathbf{G}_{j}$ since, typically, they are the ones that give more contribution to a Fourier series. With this description, $\mathbf{G}_{j}$ connects $K$ to its equivalent points in the other Brillouin zones. As indicated in Fig. 3, let us choose
@@ A $C_{3}$ symmetry operator acts on the displacement vector and the wave function as follows [54]
We notice that Eq. (16) is important because it permits us to identify reciprocal lattice vectors (Fourier modes) that contribute to the Fourier series (9) with the same strength. This identification can be done as follows: First, we take some initial $\mathbf{G}$ where we know the contribution to the Fourier series. Inserting it into (16) then permits us to find a new reciprocal lattice vector $\mathbf{G}$ that has the same associated Fourier coefficient. The same procedure can then be iterated to find additional $\mathbf{G}$ until the procedure naturally truncates. Since we already have a coefficient $w$ corresponding to $\mathbf{G}=(0,0)$ as one of our most important reciprocal lattice vectors, we can use this as the starting point. Using this procedure, we find that there are three reciprocal lattice vectors,
@@
The inter-layer tunneling can now be approximated as
@@
where the intra-layer potential $\Delta_{t,b}$ and inter-layer tunneling are given by (7) and (20), respectively. We would like to highlight that this result is identical to Ref. [22].
@@
Next, we want to find parameters $V,\psi,w,$ and $m^{*}$ corresponding to a real material. We note that they may be readily obtained by fitting the eigenvalues of Hamiltonian (21) for both the unshifted $\mathbf{d}_{0}$ and shifted $\mathbf{d}_{\pm 1}$ structures to the energy data from the DFT calculation. We stress that we only consider a small region of radius 0.04 $\pi/a_{0}$ around the $\mathbf{K}$ valley since we are only interested in the low energy region near it. We want to accurately model the region where the two topmost valence bands behave approximately quadratically. After performing the fit, we found that $(V,\psi,w,m^{*})\approx(8.66\ \text{meV},-89.97^{\circ},-8.34\ \text{meV},0.63m_{e})$ where $m_{e}$ is the mass of an electron. This result is in close agreement with the literature [22]. Our results differ only slightly from the results in the literature because we performed a full relaxation with approximately zero pressure. That is, we obtained in-plane lattice constants while considering the full bilayer systems and did not work with lattice constants obtained for the single-layer system, such as those that could be obtained from Reference [58]. While the small corrections to constants are not substantial, they build confidence in literature results and are of value for this reason. We also note that since $V$ is related to the bias between the two topmost valence bands, different choices of DFT flavors can result in different values of $V$ from what is presented here. Indeed, underestimated/overestimated energy gaps are a well-known DFT problem.
@@ II.2 Twisted case
Next, we now want to apply a relative twist between layers. Here, we imagine that we can apply the twist by rotating a layer around a Mo atom of the AA-stacked configuration ($d_{0}$ structure) - we will restrict ourselves to small angles. Such a small twist can be viewed as a collection of local displacements of atoms of the top layer with respect to the atoms of the same type in the bottom layer. Taking only the first order of the Taylor expansion for the rotation matrix, the local displacement $\mathbf{d}$ can be approximated as $\theta\hat{z}\times\mathbf{r}$.
@@
where $\Delta_{t,b}$ and $\Delta_{T}$ are defined in Eq. (7) and (20) by replacing $\mathbf{d}$ with $\theta\hat{z}\times\mathbf{r}$, which introduces a position dependence. It is important to note that position dependence is periodic and involves a larger unit cell than graphene. This larger unit cell implies a smaller Brillouin zone, which gives rise to moire bands, as we will see later.
@@ II.3 Review of band structure and topology for unpressurized twisted TMD
As a next step before we apply pressure to the twisted TMD, it is useful to review some of its properties to be able to highlight changes. The first interesting observable quantity is the band structure. To reveal the band structure of the MoTe2 moiré system, we numerically diagonalize the moiré Hamiltonian (22) using a plane-wave basis,
@@
$\begin{split}F_{xy}(\mathbf{k}_{l})\equiv\ln\bigl{[}U_{x,n}(\mathbf{k}_{l})&U_{y,n}(\mathbf{k}_{l}+\bm{x})\\ &\times U_{x,n}(\mathbf{k}_{l}+\bm{y})^{-1}U_{y,n}(\mathbf{k}_{l})^{-1}\bigr{]}.\end{split}$ (25)
@@
To better understand how valley Chern numbers change with twist angles, we have plotted the valley Chern number for the first four bands as a function of twist angle $\theta$ in Fig. 7. Additionally, Fig. 7(d) shows the gap $\Delta\epsilon_{ij}$ between bands $i$ and $j$ (the difference between the minimum energy of the $i$-th band and the maximum energy of the lower adjacent $j$-th band). From this plot, we can see that the valley Chern number of the first band remains at a value of -1 as we increase the twist angle, while the second and third bands touch and change the valley Chern number from (+1,0) to (-1,+2) at the twist angle of around 2.3°, where bands 2 and 3 touch. The band structure at the twist angle $\theta=2.3^{\circ}$ is shown in Fig. 6(b).
Figure 7: (a-c) The valley Chern number of the top three bands along various twist angles. For all considered twist angles, the valley Chern number of the first band remains as -1, while the second and third bands change valley Chern number around 2.3°. (d) The energy gap $\Delta\epsilon_{ij}$ between the $i$-th and $j$-th bands is computed as the difference between the minimum energy of the $i$-th band and the maximum energy of the $j$-th band. At around 2.3°, the gap between the second and third band closes, which is accompanied by a change in valley Chern number.

### v4

# Effective K valley Hamiltonian for TMD bilayers under pressure and application to twisted bilayers with pressure-induced topological phase transitions

@@ Abstract
Motivated by recent studies on topologically non-trivial moiré bands in twisted bilayer transition metal dichalcogenides (TMDs), we study MoTe2 bilayer systems subject to pressure, which is applied perpendicular to the material surface. We start our investigation by first considering an untwisted bilayer system with an arbitrary relative shift between layers; a symmetry analysis for this case permits us to obtain a simplified effective low-energy Hamiltonian valid near the important $\mathbf{K}$ valley region of the Brillouin zone. Ab initio density functional theory (DFT) was then employed to obtain relaxed geometric structures for pressures within the range of 0.0 - 3.5 GPa and corresponding band structures. The DFT data were then fitted to the low-energy Hamiltonian to obtain a pressure-dependent Hamiltonian. We then apply our model to a twisted system by treating the twist as a position-dependent shift between layers - here, we assume rigid layers, which is a crucial simplification. In summary, this approach allowed us to obtain the explicit analytical expressions for a Hamiltonian that describes a twisted MoTe2 bilayer under pressure. Our Hamiltonian then permitted us to study the impact of pressure on the band topology of the twisted system. As a result, we identified many pressure-induced topological phase transitions as indicated by changes in valley Chern numbers. Moreover, we found that pressure could be employed to flatten bands in some of the cases we considered.
@@ I Introduction
Exotic phenomena have also been studied in twisted TMDs (tTMDs). Experimental works, for instance, report the observation of Mott insulating state [19, 20] and quantum anomalous Hall effect [21]. On the theoretical side, twisted MoTe2 homobilayers have been predicted to exhibit nontrivial topological moiré bands [22], which indicates the existence of Chern insulators in this material. This prediction was followed by experimental results of several groups [23, 24, 25, 26, 27, 28, 29], which confirmed signatures of the theoretical predictions in their data. One might also add another TMD layer to obtain so-called trilayer TMDs. Trilayer TMDs give additional room for modification and have been shown to give rise to a plethora of Chern transitions[30]. Twisted TMDs also act as an exciting playground for simulating idealized models such as Hubbard [31, 32] and generalized Kane-Mele models [22, 33]. In some recent work [34], authors managed to construct an interacting model Hamiltonian. In the same work Ref. [34], they considered different twist angles and managed to predict transitions between various phases, including quantum anomalous Hall insulator (QAHI), quantum spin Hall insulator (QSHI), and various antiferromagnetic phases.
@@
This additional knob could be chosen in various ways. For instance, it could be circularly polarized light, which has been studied for twisted bilayer graphene [35, 36, 37, 38, 39] and TMDs [40]. Another additional knob that can be introduced to the moiré system is pressure. There are tremendous efforts underway, both experimentally [41, 42, 43, 44] and theoretically [45, 46, 47, 48], to study the effect of pressure on graphene-based moiré systems to employ it to tune its physical properties. As for TMDs, Ref. [49] studied the effect of pressure on the stability of the fractional Chern insulator phase in twisted WSe2. We note that the focus of our current work differs substantially from theirs in that we derive explicit pressure-dependent parameters for a simplified model and in that we only study non-interacting topological phases rather than the interesting interacting phase discussed in their work. Other research groups also try to introduce electric fields [50] and strains [51, 52, 53, 54] to further tune the properties of graphene-based moiré materials.
@@
In this paper, we want to help simplify the study of tTMDs subjected to pressure. We present an explicit pressure-dependent low-energy Hamiltonian for a twisted MoTe2 homobilayer in the non-interacting limit to achieve this goal. The model was derived from a general symmetry-based guess [55]. A first-principle density functional theory (DFT) approach [56, 57] was employed to obtain band structures under pressure. In this work, we will focus our attention on the bands in a small region around the $\mathbf{K}$ valley. Of course, this does not capture all the physics since there is for example contributions from the $\Gamma$ point. However, for the energy range of interest, $\mathbf{K}$ valleys contribute significantly to the density of states and will therefore play an important role in the physics. Here, band structures were obtained via DFT computations at each pressure. The results were then fitted to a low-energy Hamiltonian, which was obtained from a symmetry analysis, which allowed us to obtain explicit expression of pressure-dependent coupling parameters. We also limit the range of pressures in which our model is valid. Further investigation then leads us to the prediction of pressure-induced topological phase transitions as indicated by jumps in valley Chern numbers that correspond to specific energy bands.
@@
The remainder of the paper is structured as follows. In Section II, we present the low-energy Hamiltonian of the twisted TMD and explain the symmetry-based methods that were employed to obtain it. We demonstrate the steps to obtain the coupling parameters by fitting them to DFT data and explain the methods to compute valley Chern numbers and energy bands. Then, in Section III, we show how we use DFT computations to study structures under pressure and obtain their energy bands. Subsequently, we use this data to find the expression of the pressure-dependent coupling parameters of the Hamiltonian. We further investigate the physics of the pressurized twisted TMD by computing valley Chern number of particular energy bands within the pressure range of 0 - 3.5 GPa for twist angle $\theta$ of $1.4^{\circ}$ and $2.3^{\circ}$ and observe topological transitions that occur in our pressure range. Lastly, we conclude by summarizing our main findings and discussing the potential usage of our effective Hamiltonian in TMDs subject to additional perturbations in Section IV.
@@ II Material without pressure
This study focuses on MoTe2 homobilayer systems -including twisted bilayers. Fig. 1 illustrates such systems for the untwisted case - specifically, the MoTe2 bilayer in the AA and AB/BA stacking configurations. Before we study the twisted bilayer system, we first discuss the untwisted bilayer case to gain some physical understanding. The discussion in this section is detailed to ensure that the work is both pedagogical and self-contained.
@@
To understand the geometry of layer shifts, we consider the so-called AA stacked MoTe2 bilayer system first (Fig. 1(a) –here, atomic positions of atoms in the $x-y$ plane of one layer are exactly above those in the second layer). We may introduce a displacement vector
@@
which allows the top layer to be shifted with respect to the bottom layer. Choosing $n=\pm 1$ corresponds to the so-called AB/BA stacking (Fig. 1(b)) while $n=0$ corresponds to the AA stacking.
@@
We allowed the system to fully relax at a pressure of 0.0 GPa to obtain accurate lattice constants. With this approach, we computed the following lattice constants $a_{0}$ (see Eq. (1) and Fig. 1). For the unshifted case ($\mathbf{d}_{0}$) we found $a_{0}=3.482$ while for the shifted case($\mathbf{d}_{\pm 1}$) $a_{0}=3.487$. These numbers are very close to the lattice constants obtained for monolayer MoTe2 using DFT calculation with the LDA [59]. They differ only slightly due to coupling between the layers. Additionally, the vertical distance between the Mo atoms of each layer ($D_{z}$) was found to be 7.79Å for the unshifted structure and 6.91Å for the shifted structures, which also agrees with Ref. [22]. We note that we also performed comparisons to GGA[60] and GGA with van der Waals corrections[61] but we did not observe very significant changes (similarly for the pressure dependent case later). Hence in what follows we will stick to LDA computations.
@@
It is clear from Fig. 2 that the two topmost energy bands near the so-called $\mathbf{K}$-point behave approximately quadratically. It is worth noting that, like other TMD materials, MoTe2 features a strong spin-orbit coupling that causes the splitting of the top valence bands. Particularly, they correspond to spin-up ($\uparrow$) states for the $\mathbf{K}$ valley (spin-down ($\downarrow$) states for $\mathbf{K^{\prime}}$ valley, which is due to the time-reversal symmetry) [62, 22]. For this reason, we are now interested in finding an effective Hamiltonian that describes these two bands near $\mathbf{K}$ for arbitrary shifts between layers. Note that the model that we are developing in this paper is a simplified one and only focuses on the $\mathbf{K}$ region. As we will discuss later, contributions from other regions of the Brillouin zone - most notably the $\Gamma$ point require a separate treatment.
We switch focus to a derivation of the effective Hamiltonian for the $\mathbf{K}$ region that is based on symmetry considerations. In the cases we consider, we find that one can write a two-band model Hamiltonian as
@@
which acts on the wave function of the top and bottom layers $(\phi_{t},\phi_{b})$. (Note: One may confirm that the upper band near $\mathbf{K}$ has contributions mainly from the top layer, and the lower band has contributions mainly from the bottom layer, which suggests these identifications.)
@@ II.1.1 Intra-layer energies and their shift dependence
Figure 3: The reciprocal lattice vectors $\mathbf{G}_{j}$ connect point $\mathbf{K}$ to the nearest equivalent $\mathbf{K}$ points.
Here, $\mathbf{G}_{j}$ are the reciprocal lattice vectors that connect the origin, which is the $\mathbf{K}$ valley, to the other equivalent $\mathbf{K}$ valleys. We only choose the lowest-harmonics $\mathbf{G}_{j}$ since, typically, they are the ones that give more contribution to a Fourier series. With this description, $\mathbf{G}_{j}$ connects $\mathbf{K}$ to its equivalent points in the other Brillouin zones. As indicated in Fig. 3, let us choose
@@ A $C_{3}$ symmetry operator acts on the displacement vector and the wave function as follows [55]
We notice that Eq. (16) is important because it permits us to identify reciprocal lattice vectors (Fourier modes) that contribute to the Fourier series (9) with the same strength. This identification can be done as follows: First, we take some initial $\mathbf{G}$ for which we know the contribution to the Fourier series. Inserting it into (16) then permits us to find a new reciprocal lattice vector $\mathbf{G}$ that has the same associated Fourier coefficient. The same procedure can then be iterated to find additional $\mathbf{G}$ until the procedure naturally truncates. Since we already have a coefficient $w$ corresponding to $\mathbf{G}=(0,0)$ as one of our most important reciprocal lattice vectors, we can use this as the starting point. Using this procedure, we find that there are three reciprocal lattice vectors,
@@
The inter-layer tunneling in Equation (9) can now be approximated as
@@
where the intra-layer potential $\Delta_{t,b}$ and inter-layer tunneling are given by (7) and (20), respectively. We would like to highlight that this result is identical to that of Ref. [22].
@@
Next, we want to find parameters $V,\psi,w,$ and $m^{*}$ corresponding to a real material. We note that they may be readily obtained by fitting the eigenvalues of Hamiltonian (21) for both the unshifted $\mathbf{d}_{0}$ and shifted $\mathbf{d}_{\pm 1}$ structures to the energy data from the DFT calculation. Importantly, this fit is performed simultaneously for both configurations using a combined cost function. We stress that we only consider a small region of radius 0.04 $\pi/a_{0}$ around the $\mathbf{K}$ valley since we are only interested in the low energy region near it. We want to accurately model the region where the two topmost valence bands behave approximately quadratically. After performing the fit, we found that $(V,\psi,w,m^{*})\approx(8.66\ \text{meV},-89.97^{\circ},-8.34\ \text{meV},0.63m_{e})$ where $m_{e}$ is the electron mass. Our results differ only slightly from the results in the literature [22] because we performed a full relaxation with approximately zero pressure. That is, we obtained in-plane lattice constants while considering the full bilayer systems and did not work with lattice constants obtained for the single-layer system, such as those that could be obtained from Reference [59]. While the small corrections to constants are not substantial, they build confidence in literature results and are of value for this reason. We also note that since $V$ is related to the bias between the two topmost valence bands, different choices of DFT flavors can result in different values of $V$ from what is presented here. Indeed, underestimated/overestimated energy gaps are a well-known DFT problem.
@@ II.2 Twisted case
Next, we now want to apply a relative twist between layers. Typically in literature [55, 22] this is done by performing a rigid rotation for both layers (or one of them). A more precise treatment would incorporate relaxation effects for the moiré lattice which has been done for the twisted bilayer graphene [63, 64]. Thus, the model developed in this work offers qualitative insights for physical properties.
Here, we imagine that we can apply the twist by rotating a layer around a Mo atom of the AA-stacked configuration ($d_{0}$ structure) - we will restrict ourselves to small angles. Such a small twist can be viewed as a collection of local displacements of atoms of the top layer with respect to the atoms of the same type in the bottom layer. Taking only the first order of the Taylor expansion for the rotation matrix, the local displacement $\mathbf{d}$ can be approximated as $\theta\hat{z}\times\mathbf{r}$.
@@
where $\Delta_{t,b}$ and $\Delta_{T}$ are defined in Eq. (7) and (20) by replacing $\mathbf{d}$ with $\theta\hat{z}\times\mathbf{r}$, which introduces a position dependence. It is important to note that position dependence is periodic and involves a larger unit cell. This larger unit cell implies a smaller Brillouin zone, which gives rise to moiré bands, as we will see later.
We should stress that we have made a significant simplification for this section, namely assuming that the lattice structure is rigid under rotation. That is, we neglected the relaxation effects that occur for rotated lattices. In our discussion of the topological phases that follow, we expect that such a simplification will not qualitatively alter our results, given the robustness of topological phases. However, we caution that our model is a simplified one that also does not include interaction effects, and it cannot, therefore, be expected to capture correlated phases such as a fractional Chern insulator phase, like the one discussed in [29].
@@ II.3 Review of band structure and topology for unpressurized twisted TMD
As a next step before we apply pressure to the twisted TMD, it is useful to review some of its properties to be able to highlight changes. Including those details is useful to keep our work self-contained. The first interesting observable quantity is the band structure. To reveal the band structure of the MoTe2 moiré system, we numerically diagonalize the moiré Hamiltonian (22) using a plane-wave basis,
@@
$\begin{split}F_{xy}(\mathbf{k}_{l})\equiv&\ln\bigl{[}U_{x,n}(\mathbf{k}_{l})U_{y,n}(\mathbf{k}_{l}+\bm{x})\\ &\times U_{x,n}(\mathbf{k}_{l}+\bm{y})^{-1}U_{y,n}(\mathbf{k}_{l})^{-1}\bigr{]}.\end{split}$ (25)
@@
To better understand how valley Chern numbers change with twist angles, we have plotted the valley Chern number for the first three bands as a function of twist angle $\theta$ in Fig. 7. Additionally, Fig. 7(d) shows the gap $\Delta\epsilon_{ij}$ between bands $i$ and $j$ (the difference between the minimum energy of the $i$-th band and the maximum energy of the lower adjacent $j$-th band). From this plot, we can see that the valley Chern number of the first band remains at a value of -1 as we increase the twist angle, while the second and third bands touch and change the valley Chern number from (+1,0) to (-1,+2) at a twist angle of around 2.3°, where bands 2 and 3 touch. The band structure at the twist angle $\theta=2.3^{\circ}$ is shown in Fig. 6(b).
Figure 7: (a-c) The valley Chern number of the top three bands along various twist angles. For all considered twist angles, the valley Chern number of the first band remains -1, while the second and third bands change valley Chern number around 2.3°. (d) The energy gap $\Delta\epsilon_{ij}$ between the $i$-th and $j$-th bands is computed as the difference between the minimum energy of the $i$-th band and the maximum energy of the $j$-th band. At around 2.3°, the gap between the second and third band closes, which is accompanied by a change in valley Chern number.